OpenStudy (anonymous):

Burton Bush wants to retire in Arizona when he is 80 years of age. Burton, who is now 55, believes he will need $400,000 to retire comfortably. To date, he has set aside no retirement money. If he gets an interest of 6 percent compounded annually, he will have to invest today: (Use the tables in the handbook) $92,300 $69,900 $96,500 $93,200 None of these

hero (hero):

Hint: \[A = P(1 + \frac{r}{n})^{nt}\] A = 400,000 r = .06 t = 12 n = 80 - 55 P = one of the answer choices